# Gamma-Ray Burst Light Curve Reconstruction: A Comparative Machine and Deep Learning Analysis

### v4

@@ 2.3.8 Kolmogorov-Arnold Networks
The KAN (Kolmogorov-Arnold Network) model was first introduced by Liu et al. (2024) and has since been studied extensively. One of its key advantages is its ability to recognize complex patterns and relationships in data.
@@ 2.4.1 Fourier Transform
This concept was implemented to reconstruct GRB LCs by transforming time-domain flux values into frequency-domain coefficients. The initial data was first processed using the Fast Fourier Transform (FFT) (Oberst, 2007) on the flux values. The interpolation technique imputes the new flux values, and the uncertainties were predicted using the normal distribution. This process helps us understand how flux values may vary over time, taking uncertainties into account. Fig. 3h is the FFT synthesis flow diagram that shows how the FFT works. It shows how the FFT combines two 4-point spectra into an 8-point spectrum. The ”×$\times S$” symbol represents the signal multiplied by a sinusoid of appropriate frequency.
@@ 2.5 SARIMAX-based Kalman Smoothing Model
The SARIMAX model captures the observed flux data’s temporal dependencies, trends, and noise. An ARIMA model (Ho & Xie, 1998) of order (p,d,q) is tuned accordingly to model the LC’s temporal dynamics. Here, p is the number of autoregressive terms, d is the number of non-seasonal differences, and q is the number of lagged forecast errors in the prediction equation. An order of (1,1,1) was used that often captures basic temporal structure with fewer parameters, avoiding overfitting while reducing autocorrelation in residuals. In our model, we applied Kalman smoothing to refine the SARIMAX output. Kalman smoothing works in such a way that it estimates the system’s state using measurements from both past and future time steps. Let $\mathbf{x}_{t}$ denote the state vector at time $t$, modeled as:
@@ 4 Results
Decreasing the uncertainty associated with the W07 parameters is an objective of LCR. To evaluate this, the error fractions, symbolized by $EF$, are calculated for every model parameter in both the primary datasets and after reconstructing them using the model. The error fractions for the three W07 parameters as described in Dainotti et al. 2023b are represented in Eq. 41, 42, and 43:
@@
%$\%_{DEC}=\frac{\left|EF^{\rm{after}}_{X}\right|-\left|EF^{\rm{before}}_{X}\right|}{\left|EF^{\rm{before}}_{X}\right|}\times 100.$ (44)
@@ 4.1 Results from the Willingale Model
The reconstruction results for all four classes of GRBs under 10% and 20% noise levels are illustrated in Fig. 4. For all classes of GRBs, the LCs are reconstructed to demonstrate how increasing noise levels affect the distribution of data points near the W07 fit. As anticipated, increasing the noise level generally leads to higher uncertainty due to the broader uncertainty around the best-fit line. However, in our case, the decrease in uncertainty does not follow this pattern with increasing noise levels. This is because we are considering the entire sample. The data points deviate significantly from the model function for GRBs exhibiting flares and breaks. As a result, with more significant noise levels, the decrease in uncertainty is less pronounced. Additionally, the EFs for the W07 model applied to 545 good GRBs at both 10% and 20% noise levels are summarized in Table LABEL:tab:ALL_Table. The results include the histogram distribution of the percentage decrease for all three W07 parameters at these noise levels.
A reduction of 24.5% for $\log T_{a}$, 25.7% for $\log F_{a}$, and 36.2% for $\alpha$ is observed at a 10% noise level. However, for the subset containing only 218 good GRBs, it performs the best showcasing 33.3% for $\log T_{a}$, 35.0% for $\log F_{a}$and 43.3% for $\alpha$ with no additional outliers. As expected, the reductions were marginally smaller, with 21.2% for $\log T_{a}$, 22.9% for $\log F_{a}$, and 34.7% for $\alpha$, for a 20% noise level.
@@ 4.2 Results from the ML and statistical methods
Figure 6: Reconstruction of LCs for all four varieties of GRBs are shown in a grid with four types of GRBs (left to right): i) Good GRBs (Column 1); ii) a GRB LC with a break towards the end (Column 2); iii) flares or bumps in the afterglow (Column 3); iv) flares or bumps with a double break towards the end of the LC (Column 4) and the models (top to bottom): i) Bi-LSTM Model(Row 1); ii) SARIMA-based Kalman Model (Row 2); iii) CGAN (Row 3); iv) KAN model (Row 4); v) Attention UNet model (Row 5).
@@
Figure 8: Distribution plot of all three W07 parameters in a grid with parameters (left to right): i) $\log F_{a}$ (Column 1) ii) $\log T_{a}$ (Column 2) iii) $\alpha$ (Column 3) and the models (top to bottom): i) GP-RF (Row 1); ii) Bi-LSTM (Row 2); iii) CGAN (Row 3); iv) SARIMAX-based Kalman (Row 4); v) KAN (Row 5); vi) U-Net Attention (Row 6).
@@
GRB ID $EF_{\log_{10}(T_{i})}$ $EF_{\log_{10}(F_{i})}$ $EF_{\alpha_{i}}$ $EF_{\log_{10}(T_{i})}$ RC $EF_{\log_{10}(F_{i})}$ RC $EF_{\alpha_{i}}$ RC %$\%\log_{10}(T_{i})$$\%\log_{10}(F_{i})$$\%\alpha_{i}$
@@
% $\log T_{a}$ $\log F_{a}$ $\alpha$ $\log T_{a}$ $\log F_{a}$ $\alpha$ Train MSE ($10^{-1}$) Test MSE ($10^{-1}$) $\downarrow$
@@
MLP -25.9 -28.6 -37.7 3.60 3.30 0.900 0.227 0.275
Fourier -28.8 -27.8 -36.7 3.30 3.50 0.500 0.0270 0.339
CGAN -18.8 -19.1 -24.3 3.60 4.40 3.30 0.260 0.429
Bi-LSTM -28.4 -28.2 -36.0 3.80 3.60 1.20 0.231 0.532
SARIMAX-based Kalman -19.3 -24.6 -28.9 4.10 3.70 0.900 0.814 0.825
Bi-Mamba -33.3 -33.6 -41.9 2.70 2.70 0.700 0.151 1.30
GP-RF -25.0 -27.5 -34.7 5.60 5.30 2.30 0.0551 1.33
Attention U-Net -27.5 -28.3 -38.1 3.10 3.10 0.300 0.206 1.34
KAN model -15.5 -25.4 -27.0 7.70 3.50 1.60 0.423 1.74
GP (W07) -22.8 -23.2 -33.5 3.10 3.30 1.10 8.56 3.63
W07 model (10%) -24.5 -25.7 -36.2 2.40 2.40 0.900 - -
W07 model (20%) -21.2 -22.9 -34.7 2.80 2.80 1.50 - -
218 Good GRBs Dainotti et al. (2023b)
SARIMAX-based Kalman -21.4 -26.3 -30.4 2.20 1.80 0.400 0.191 0.192
MLP -24.1 -27.8 -38.9 0.900 0.400 0.400 0.230 0.268
CGAN -18.3 -20.7 -26.5 1.80 2.30 2.30 0.230 0.383
Fourier -26.3 -28.2 -38.2 0.900 0.900 0 0.0900 0.480
Bi-LSTM -26.1 -28.0 -37.3 1.30 1.80 1.30 0.210 0.626
Bi-Mamba -32.7 -34.3 -43.6 0.400 0.400 0.400 0.110 0.850
GP-RF -24.2 -27.8 -37.7 1.30 1.30 0.400 0.0468 0.920
Attention U-Net -26.5 -28.2 -38.1 0.900 0.900 0 0.173 1.07
KAN model -14.6 -25.6 -25.7 5.90 1.30 0.400 0.336 1.61
GP (W07) -25.6 -27.9 -41.6 0.400 0 0 7.66 1.81
W07 model (10%) -33.3 -35.0 -43.3 0 0 0 - -
W07 model (20%) -29.5 -31.2 -40.6 0 0 0 - -
Table 5: Summary of average decrease in the % uncertainty along with the % increase in outliers across different reconstruction processes and their parameters. The train MSE and test MSE for all the models across 5 K-Fold CV are provided. The models are sorted in ascending order based on Test MSE values (top to bottom). GP, W07 (10%) and W07 (20%) are shown separately at the bottom for both 545 GRBs and 218 Good GRBs.
@@ 5 Summary and Conclusion
  • The W07 model (10% noise level) is most effective for reconstructing 218 Good GRBs, achieving the highest uncertainty reduction across all parameters while eliminating outliers. However, its performance diminishes with flares or breaks in the data, limiting its applicability. The model fundamentally relies on the Willingale function, which is designed to represent a generalized LC structure. As such, it does not explicitly account for complex features like flares or breaks often observed in half of the GRBs. Similarly, the GP approach shows reduced reliability in handling GRBs with flares or breaks. Although it provides a more flexible non-parametric framework compared to the fixed functional form of W07, it still tends to follow the overall smooth trend of the LC and has issues in reconstructing transient deviations accurately.
@@
  • The MLP model provides the lowest 5-Fold Test MSE of 0.0275 among deep learning models, making it the best option when prioritizing reconstruction accuracy. It also offers moderate uncertainty reduction (25.9% for $\log T_{a}$, 28.6% for $\log F_{a}$, and 37.7% for $\alpha$) while maintaining low outlier rates across all parameters.
@@
  • The Bi-Mamba model achieves a higher test MSE of 0.130 but excels in minimizing outliers, particularly for $\log T_{a}$ and $\log F_{a}$. It also exhibits the highest uncertainty reduction across all parameters (33.3% for $\log T_{a}$, 33.6% for $\log F_{a}$, and 41.9% for $\alpha$).
  • The Bi-LSTM model delivers comparable performance to MLP with a test MSE of 0.0532, almost twice that of MLP. It is an effective alternative, especially when reducing computational complexity is a priority.
@@
  • For $\log T_{a}$, Bi-Mamba reduces the outliers by 25.0% compared to MLP and 28.9% compared to Bi-LSTM. Similarly, for $\log F_{a}$, Bi-Mamba achieves a reduction of 18.8% compared to MLP and 22.8% compared to Bi-LSTM. For $alpha$, Bi-MAMBA reduces the outliers by 22.2% compared to MLP and 41.7% compared to Bi-LSTM.
@@
  • The Fourier model, CGAN, and SARIMAX-based Kalman model compared to the other models, demonstrate lower Test MSE values (ranging from 0.0339 to 0.0825) but are not the best overall choices due to trade-offs in handling uncertainty, outliers, and model overfitting (as indicated by the discrepancy in the Train and Test losses).
@@
  • The Attention U-Net, GP-RF, and KAN models show higher test MSE and poor outlier control than the other tested models, making them less suitable for this specific reconstruction task.
    We conclude that the MLP model is the most effective approach for GRB reconstruction, achieving the lowest Test MSE. While the Bi-Mamba exhibit lower performance in test MSE, it is particularly effective in minimizing outliers and reducing the overall uncertainty in parameters.
@@
In comparison to the prior study by Dainotti et al. (2023b), we have the following results:
@@
  • As highlighted by Dainotti et al. (2023b), a study of 218 Good GRBs revealed an average reduction in uncertainties of 37% across all parameters with a 10% noise level and 34% at a 20% noise level on modeling with the W07 model. In contrast, the GP model exhibited a slightly lower average reduction of 31%.
@@
  • In our work, we include this exact dataset of 218 GRBs for a fair comparison (see Table 5). The Bi-MAMBA model achieves a percentage decrease in uncertainty for the $\log F_{a}$ and $\log T_{a}$ parameters that is nearly comparable to that of the W07 method (with 10% noise), with only an average difference of 1.9% in favor of W07. However, for the parameter $\alpha$, Bi-MAMBA slightly outperforms W07, achieving a 43.6% reduction in uncertainty compared to 43.3% with W07, indicating a difference of 0.7% but a noteworthy improvement. Compared to the GP model, Bi-MAMBA outperforms for all three parameters. Additionally, the 5-fold CV results show that both training and testing MSE values for Bi-MAMBA are lower than those of GP, further demonstrating its reliability. Another advantage of Bi-MAMBA is that it is model independent while achieving performance equivalent to W07.
@@
  • This study effectively reduces parameter uncertainty for the subset of 218 well-behaved GRBs. However, when the approach is extended to the full dataset encompassing all four GRB categories, the GP method exhibits reduced performance, as outlined in Dainotti et al. (2023b). In contrast, the other ML and DL models, particularly Bi-MAMBA, maintain a similar level of precision in reducing parameter uncertainty. This highlights the strength of our models in reconstructing more complex GRB LCs that include features such as flares or breaks. Furthermore, both training and testing MSE values remain low, demonstrating the continued reliability of the models. Although the number of outliers increases, from an average of 0.13% for the 218 GRBs to 2.5% across 545 GRBs, this is expected. This rise is expected as the Willingale function is not designed to model the behavior of more complex LCs.
@@
  • In summary, this work undertakes a more comprehensive and challenging analysis than Dainotti et al. (2023b) by applying advanced ML and DL methods to a significantly larger and more diverse GRB dataset. We provide more insight on the strengths and weaknesses of each approach by methodically evaluating a larger variety of GRBs, including those with intricate characteristics like flares and breaks. These advancements not only extend the scope of previous studies but also highlight the reliability and generalizability of our models.
@@
The substantial decrease in uncertainties attained through these reconstructions enhances the applicability of the GRB plateau parameters for cosmological tools (Dainotti et al., 2022g, c, 2023a).
@@
Among the different evaluation aspects, the most important parameters in our analysis are $\log T_{a}$, $\log F_{a}$, and $\alpha$, as they carry physical significance: $\alpha$ is crucial for theoretical modeling, as it describes the post-plateau decay behavior. This is relevant for the testing of the standard fireball model (Piran, 1999), where the reduction of the $\alpha$ parameter is crucial to obtain reduced uncertainties on the closure relationship (Srinivasaragavan et al., 2020b; Dainotti et al., 2021, 2024a). It refers to the relationship between the spectral and temporal indices of a GRB’s afterglow. $\log T_{a}$ and $\log F_{a}$ are key for identifying empirical correlations, such as the Dainotti relations in two and three dimensions (Dainotti et al., 2008, 2010, 2011, 2013, 2016, 2020a, 2020b, 2022b; Levine et al., 2022). The 2D Dainotti relation relates the end time of the plateau and its luminosity, while the three-dimensional relation adds to the 2D relation the peak of the prompt emission parameter. These relations are of particular interest in cosmological applications and also for distinguishing among theoretical models.
@@
Therefore, we consider these three parameters as the most important outputs of our reconstruction. Among all models tested, Bi-MAMBA demonstrates the most significant reduction in error for the three parameters. Thus, we recommend it as the preferred model when accurate estimation of these physically meaningful quantities is the main objective. The MSE is the second most important metric in our study. While it does not capture physical significance directly, it quantifies the overall performance and accuracy of the ML models, making it essential for model comparison. The MLP model has the lowest Test MSE. Outlier analysis is complementary and helps interpret cases where reconstruction might be less representative or behave anomalously.
@@
This reconstruction framework has been developed exclusively using Swift LCs so far; however, we intend to extend its application to additional current missions such as SVOM (Atteia et al., 2022) and Einstein Probe (Yuan et al., 2022), as well as future missions, including THESEUS (Amati et al., 2018) and HiZ-GUNDAM (Yonetoku et al., 2024). Our extension will also incorporate data across multiple wavelengths. Expanding this work to optical wavelengths is especially timely, given the availability of the most extensive optical catalog to date (Dainotti et al., 2020a, 2022b, 2024a, 2024c).

### v5

@@ 2.3.8 Kolmogorov-Arnold Networks
The KAN (Kolmogorov-Arnold Network) model was first introduced by liu2024kan and has since been studied extensively. One of its key advantages is its ability to recognize complex patterns and relationships in data.
@@ 2.4.1 Fourier Transform
This concept was implemented to reconstruct GRB LCs by transforming time-domain flux values into frequency-domain coefficients. The initial data was first processed using the Fast Fourier Transform (FFT) (OberstUlrich) on the flux values. The interpolation technique imputes the new flux values, and the uncertainties were predicted using the normal distribution. This process helps us understand how flux values may vary over time, taking uncertainties into account. Fig. 3h is the FFT synthesis flow diagram that shows how the FFT works. It shows how the FFT combines two 4-point spectra into an 8-point spectrum. The $\times S$” symbol represents the signal multiplied by a sinusoid of appropriate frequency.
@@ 2.5 SARIMAX-based Kalman Smoothing Model
The SARIMAX model captures the observed flux data’s temporal dependencies, trends, and noise. An ARIMA model (HO1998213) of order (p,d,q) is tuned accordingly to model the LC’s temporal dynamics. Here, p is the number of autoregressive terms, d is the number of non-seasonal differences, and q is the number of lagged forecast errors in the prediction equation. An order of (1,1,1) was used that often captures basic temporal structure with fewer parameters, avoiding overfitting while reducing autocorrelation in residuals. In our model, we applied Kalman smoothing to refine the SARIMAX output. Kalman smoothing works in such a way that it estimates the system’s state using measurements from both past and future time steps. Let $\mathbf{x}_{t}$ denote the state vector at time $t$, modeled as:
@@ 4 Results
Decreasing the uncertainty associated with the W07 parameters is an objective of LCR. To evaluate this, the error fractions, symbolized by $EF$, are calculated for every model parameter in both the primary datasets and after reconstructing them using the model. The error fractions for the three W07 parameters as described in dainotti2023stochastic are represented in Eq. 41, 42, and 43:
@@
$\%_{DEC}=\frac{\left|EF^{\rm{after}}_{X}\right|-\left|EF^{\rm{before}}_{X}\right|}{\left|EF^{\rm{before}}_{X}\right|}\times 100.$ (44)
@@ 4.1 Results from the Willingale Model
The reconstruction results for all four classes of GRBs under 10% and 20% noise levels are illustrated in Fig. 4. For all classes of GRBs, the LCs are reconstructed to demonstrate how increasing noise levels affect the distribution of data points near the W07 fit. As anticipated, increasing the noise level generally leads to higher uncertainty due to the broader uncertainty around the best-fit line. However, in our case, the decrease in uncertainty does not follow this pattern with increasing noise levels. This is because we are considering the entire sample. The data points deviate significantly from the model function for GRBs exhibiting flares and breaks. As a result, with more significant noise levels, the decrease in uncertainty is less pronounced. Additionally, the EFs for the W07 model applied to 521 good GRBs at both 10% and 20% noise levels are summarized in Table LABEL:tab:ALL_Table. The results include the histogram distribution of the percentage decrease for all three W07 parameters at these noise levels.
A reduction of 24.5% for $\log T_{a}$, 25.7% for $\log F_{a}$, and 36.2% for $\alpha$ is observed at a 10% noise level. However, for the subset containing only 207 good GRBs, it performs the best showcasing 33.3% for $\log T_{a}$, 35.0% for $\log F_{a}$% and 43.3% for $\alpha$ with no additional outliers. As expected, the reductions were marginally smaller, with 21.2% for $\log T_{a}$, 22.9% for $\log F_{a}$, and 34.7% for $\alpha$, for a 20% noise level.
@@ 4.2 Results from the ML and statistical methods
Figure 6: Reconstruction of LCs for all four varieties of GRBs are shown in a grid with four types of GRBs (left to right): i) Good GRBs (Column 1); ii) a GRB LC with a break towards the end (Column 2); iii) flares or bumps in the afterglow (Column 3); iv) flares or bumps with a double break towards the end of the LC (Column 4) and the models (top to bottom): i) Bi-LSTM Model(Row 1); ii) SARIMA-based Kalman Model (Row 2); iii) CGAN (Row 3); iv) KAN model (Row 4); v) Attention U-Net model (Row 5).
@@
Figure 8: Distribution plot of all three W07 parameters in a grid with parameters (left to right): i) $\log F_{a}$ (Column 1) ii) $\log T_{a}$ (Column 2) iii) $\alpha$ (Column 3) and the models (top to bottom): i) GP-RF (Row 1); ii) Bi-LSTM (Row 2); iii) CGAN (Row 3); iv) SARIMAX-based Kalman (Row 4); v) KAN (Row 5); vi) Attention U-Net (Row 6).
@@
GRB ID $EF_{\log_{10}(T_{i})}$ $EF_{\log_{10}(F_{i})}$ $EF_{\alpha_{i}}$ $EF_{\log_{10}(T_{i})}$ RC $EF_{\log_{10}(F_{i})}$ RC $EF_{\alpha_{i}}$ RC $\%\log_{10}(T_{i})$ $\%\log_{10}(F_{i})$ $\%\alpha_{i}$
@@
% $\log T_{a}$ % $\log F_{a}$ % $\alpha$ % $\log T_{a}$ % $\log F_{a}$ % $\alpha$ Train MSE ($10^{-1}$) Test MSE ($10^{-1}$) $\downarrow$
@@
MLP -37.2 -38.0 -41.2 1.73 2.30 1.34 0.227 0.275
Fourier -14.9 -15.2 -20.6 4.61 4.22 2.50 0.0270 0.339
CGAN -14.5 -15.4 -20.0 3.26 3.45 1.34 0.260 0.429
Bi-LSTM -21.2 -21.5 -26.1 2.50 2.88 1.34 0.231 0.532
SARIMAX-based Kalman -9.70 -13.3 -17.3 3.26 3.84 1.34 0.814 0.825
Bi-Mamba -20.8 -21.2 -27.6 2.88 2.30 1.15 0.151 1.30
GP-RF -16.3 -19.5 -23.2 4.03 4.03 2.30 0.0551 1.33
Attention U-Net -37.9 -38.5 -41.4 1.73 2.50 1.34 0.206 1.34
KAN model -4.74 -14.2 -11.6 4.99 1.92 1.73 0.423 1.74
GP (W07) -16.9 -18.6 -24.3 3.07 3.45 1.54 8.56 3.63
W07 model (10%) -18.0 -19.1 -25.2 2.30 2.30 2.11 - -
W07 model (20%) -15.8 -17.8 -23.8 2.30 2.69 2.30 - -
207 Good GRBs dainotti2023stochastic
SARIMAX-based Kalman -10.9 -15.9 -19.8 0.966 1.45 0 0.191 0.192
MLP -38.1 -39.3 -43.9 0.483 0.483 0 0.230 0.268
CGAN -16.3 -17.9 -24.6 0.966 0.966 0 0.230 0.383
Fourier -15.5 -17.3 -24.1 1.45 1.45 0.483 0.0900 0.480
Bi-LSTM -22.7 -24.0 -30.1 0.483 0.483 0 0.210 0.626
Bi-Mamba -22.5 -24.4 -31.2 0.483 0.483 0 0.110 0.850
GP-RF -17.6 -21.8 -27.8 0.483 0.966 0.483 0.0468 0.920
Attention U-Net -38.8 -40.3 -44.0 0.483 0.483 0 0.173 1.07
KAN model -7.06 -17.3 -12.8 3.38 0 0 0.336 1.61
GP (W07) -17.2 -20.0 -28.8 1.93 1.45 0 7.66 1.81
W07 model (10%) -23.0 -24.9 -30.8 0.966 0.966 0 - -
W07 model (20%) -21.2 -23.2 -28.9 0 0 0 - -
Table 5: Summary of average decrease in the % uncertainty along with the % increase in outliers across different reconstruction processes and their parameters. The train MSE and test MSE for all the models across 5 K-Fold CV are provided. The models are sorted in ascending order based on Test MSE values (top to bottom). GP, W07 (10%) and W07 (20%) are shown separately at the bottom for both 521 GRBs and 207 Good GRBs.
@@ 5 Summary and Conclusion
  • The W07 model (10% noise level) is most effective for reconstructing 207 Good GRBs, achieving the highest uncertainty reduction across all parameters while eliminating outliers. However, its performance diminishes with flares or breaks in the data, limiting its applicability. The model fundamentally relies on the Willingale function, which is designed to represent a generalized LC structure. As such, it does not explicitly account for complex features like flares or breaks often observed in half of the GRBs. Similarly, the GP approach shows reduced reliability in handling GRBs with flares or breaks. Although it provides a more flexible non-parametric framework compared to the fixed functional form of W07, it still tends to follow the overall smooth trend of the LC and has issues in reconstructing transient deviations accurately.
@@
  • The MLP model provides the lowest 5-Fold Test MSE of 0.0275 among deep learning models, making it the best option when prioritizing reconstruction accuracy. It also offers a high uncertainty reduction (37.2% for $\log T_{a}$, 38.0% for $\log F_{a}$, and 41.2% for $\alpha$) while maintaining low outlier rates across all parameters.
@@
  • The Attention U-Net model achieves a higher test MSE of 0.134 but excels in minimizing outliers, particularly for $\log T_{a}$ and $\log F_{a}$. It also exhibits the highest uncertainty reduction across all parameters (37.9% for $\log T_{a}$, 38.5% for $\log F_{a}$, and 41.4% for $\alpha$).
  • In comparison, the Bi-LSTM and Bi-Mamba models show comparable levels of uncertainty reduction across the three parameters: 21.2%, 21.5%, and 26.1% for Bi-LSTM, and 20.8%, 21.2%, and 27.6% for Bi-Mamba, respectively. They are an effective alternative, especially when reducing computational complexity is a priority. Bi-Mamba also exhibits the lowest outliers for $\alpha$ of 1.15.
@@
  • For $\log T_{a}$, the Attention U-Net reduces the outliers by 1.88% compared to the MLP baseline. Similarly, for $\log F_{a}$, it achieves a 1.32% reduction, while for $\alpha$, the reduction amounts to 0.49%.
@@
  • The Fourier model, CGAN, and SARIMAX-based Kalman model, compared to the other models, demonstrate lower Test MSE values (ranging from 0.0339 to 0.0825) but are not the best overall choices due to trade-offs in handling uncertainty, outliers, and model overfitting (as indicated by the discrepancy in the Train and Test losses).
@@
  • The GP-RF and KAN models show higher test MSE and poor outlier control than the other tested models, making them less suitable for this specific reconstruction task.
    In summary, while Attention U-Net achieves the highest reduction in outliers and predictive uncertainty, its test MSE (0.134) is higher than that of the MLP. The MLP, in contrast, delivers the lowest test MSE and remains highly competitive in uncertainty reduction, with performance very close to the Attention U-Net (within 0.5–1.9% across all parameters). This balance between accuracy and stability suggests that the MLP provides a more reliable baseline for GRB LCR, while the Attention U-Net offers advantages in minimizing extreme deviations.
@@
In comparison to the prior study by dainotti2023stochastic, we have the following results:
@@
  • As highlighted by dainotti2023stochastic, a study of 207 Good GRBs revealed an average reduction in uncertainties of 26% across all parameters with a 10% noise level and 24% at a 20% noise level on modeling with the W07 model.
@@
  • In our work, we include this exact dataset of 207 GRBs for a fair comparison (see Table 5). When compared to the W07 model, which achieves uncertainty reductions of 23.0% for $\log T_{a}$, 24.9% for $\log F_{a}$, and 30.8% for $\alpha$, the Attention U-Net improves these values substantially to 38.8%, 40.3%, and 44.0%, respectively. This corresponds to relative improvements of 68.7%, 61.9%, and 42.9%, indicating that the Attention U-Net provides a more effective reduction of predictive uncertainty compared to the W07 model. Compared to the GP model, Attention U-Net outperforms for all three parameters. Additionally, the 5-fold CV results show that both training and testing MSE values for Attention U-Net are lower than those of GP, further demonstrating its reliability. Another advantage of Attention U-Net is that it is model-independent while achieving performance equivalent to W07.
@@
  • This study effectively reduces parameter uncertainty for the subset of 207 well-behaved GRBs. However, when the approach is extended to the full dataset encompassing all four GRB categories, the GP method exhibits reduced performance, as outlined in dainotti2023stochastic. In contrast, the other ML and DL models, particularly Attention U-Net, maintain a similar level of precision in reducing parameter uncertainty. This highlights the strength of our models in reconstructing more complex GRB LCs that include features such as flares or breaks. Furthermore, both training and testing MSE values remain low, demonstrating the continued reliability of the models. Although the number of outliers increases, from an average of 0.13% for the 207 GRBs to 2.5% across 521 GRBs, this is expected. This rise is expected as the Willingale function is not designed to model the behavior of more complex LCs.
@@
  • In summary, this work undertakes a more comprehensive and challenging analysis than dainotti2023stochastic by applying advanced ML and DL methods to a significantly larger and more diverse GRB dataset. We provide more insight on the strengths and weaknesses of each approach by methodically evaluating a larger variety of GRBs, including those with intricate characteristics like flares and breaks. These advancements not only extend the scope of previous studies but also highlight the reliability and generalizability of our models.
@@
The substantial decrease in uncertainties attained through these reconstructions enhances the applicability of the GRB plateau parameters for cosmological tools (dainotti2022g; dainotti2022c; dainotti2023b).
@@
Among the different evaluation aspects, the most important parameters in our analysis are $\log T_{a}$, $\log F_{a}$, and $\alpha$, as they carry physical significance: $\alpha$ is crucial for theoretical modeling, as it describes the post-plateau decay behavior. This is relevant for the testing of the standard fireball model (piran1999gamma), where the reduction of the $\alpha$ parameter is crucial to obtain reduced uncertainties on the closure relationship (closure; dainotti2021closure; dainotti2024analysis). It refers to the relationship between the spectral and temporal indices of a GRB’s afterglow. $\log T_{a}$ and $\log F_{a}$ are key for identifying empirical correlations, such as the Dainotti relations in two and three dimensions (Dainotti2008; dainotti2010a; dainotti2011a; dainotti2013determination; dainotti2016fundamental; dainotti2020a; dainotti2020b; dainotti2022b; levine2022). The 2D Dainotti relation relates the end time of the plateau and its luminosity, while the three-dimensional relation adds to the 2D relation the peak of the prompt emission parameter. These relations are of particular interest in cosmological applications and also for distinguishing among theoretical models.
@@
Therefore, we consider these three parameters as the most important outputs of our reconstruction. Among all models tested, Attention U-Net demonstrates the most significant reduction in error for the three parameters. Thus, we recommend it as the preferred model when accurate estimation of these physically meaningful quantities is the main objective. The MSE is the second most important metric in our study. While it does not directly capture physical significance, it quantifies the overall performance and accuracy of the ML models, making it essential for model comparison. By this measure, the MLP performs best, delivering the lowest test MSE while still achieving uncertainty reductions that are very close to those of the Attention U-Net. Outlier analysis is complementary and helps interpret cases where reconstruction might be less representative or behave anomalously.
@@
This reconstruction framework has been developed exclusively using Swift LCs so far; however, we intend to extend its application to additional current missions such as SVOM (Atteia et al., 2022) and Einstein Probe (yuan2022einstein), as well as future missions, including THESEUS (Amati et al., 2018) and HiZ-GUNDAM (yonetoku2024high). Our extension will also incorporate data across multiple wavelengths. Expanding this work to optical wavelengths is especially timely, given the availability of the most extensive optical catalog to date (dainotti2020optical; dainotti2022b; dainotti2024analysis; dainotti2024optical).